# Generalized local jocabians and commutative group stacks

### v2

@@ 3.2. Grothendieck’s construction
[…]
We thus know that $\psi_{\mathbb{G}_{a}}^{i}$ is an isomorphism for all $i$. It remains to treat the case of a general commutative unipotent group $G$. Any such group possesses a decomposition series whose layers $G_{\alpha}$ can be realized in a short exact sequence $0\to G_{\alpha}\to\mathbb{G}_{a}\to\mathbb{G}_{a}$. The associated long exact sequences and the lemma easily implies that $\psi_{G}^{i}$ is also an isomorphism for all $i$ (note that the domains and codomains of $\psi_{G}^{i}$ can be non-zero for $i=1$). $\Box$
@@
    The previous proposition provides a close relation between our algebraic homology $\mathbb{H}^{alg}(X)$ and the original Grothendieck’s construction $J_{*}(X)$ of. The former is a truncated version of the later, but both can be used in order to express the fppf cohomology of $X$ with coefficients in a commutative group scheme.
@@
    The extra conditions on $G$, being affine for $i>0$ and unipotent for $i>1$, does not appear in Grothendieck’s original construction. This is because his construction is originaly made using the Zariski topology instead othe fppf topology, and higher Zariski cohomology groups with coefficients in group schemes vanish outside of the unipotent context.
@@ 4. Partial generalization to more general schemes
In this section we partially extend the results of the previous sections to the more general case of schemes over arbitrary bases. For this, we restrict ourselves to the unipotent part of algebraic homology and use the formalism of commutative groups in affine stacks (in the sense of [Toe06]) as replacement for (pro) commutative group stacks. On the way, we develop the general framework of affine homology, the natural homology theory associated to affine stacks.
@@ 4.1. Affine homology
When $C=\mathbf{St}_{k}$ is the $\infty$-category of stacks over $k$, $\mathbf{Ab}C$ is naturally equivalent to $\mathbf{AbSt}_{k}$ (already used previously in this work, but in the setting of the fppf topology), and is identified with the $\infty$-category of connective complexes of sheaves of abelian groups on the site of affine $k$-schemes with the fppf topology.
@@
The $\infty$-category of commutative affine group stacks over $k$ is defined to be the full sub-$\infty$-category of $\mathbf{AbSt}_{k}$ consisting of all commutative group stacks $E$ such that for all $n\geq 0$ the underlying stack $|E[n]|\in\mathbf{St}_{k}$ is an affine stack in the sense of [Toe06]. The $\infty$-category of all commutative affine group stacks will be denoted by $\mathbf{AbChAff}_{k}$.
We warn the reader that the notation $\mathbf{AbChAff}_{k}$ can be misleading. Indeed, the $\infty$-category $\mathbf{ChAff}_{k}$ defines another $\infty$-category $Ab(\mathbf{ChAff}_{k})$, of abelian group objects in $\mathbf{ChAff}_{k}$. Recall that this is formally defined as the full sub-$\infty$-category of $Fun^{\infty}(T^{op},\mathbf{ChAff}_{k})$, consisting of all $\infty$-functors sending finite sums in $T$ to finite products in $C$, and where $T$ is the algebraic theory of abelian group (i.e. the opposite of the category of free abelian group of finite rank). The $\infty$-category $\mathbf{AbChAff}_{k}$ is smaller than $Ab(\mathbf{ChAff}_{k})$. Indeed, any affine commutative group scheme $G$ defines an object in $Ab(\mathbf{ChAff}_{k})$. However, we will see below (see proposition 4.2) that $\mathbb{G}_{m}$ is not an object in $\mathbf{AbChAff}_{k}$. We thus have canonical fully faithful $\infty$-functors $\mathbf{AbChAff}_{k}\hookrightarrow Ab(\mathbf{ChAff}_{k})\hookrightarrow\mathbf{AbSt}_{k}$, but these three $\infty$-categories are all distinct.
@@
An important first observation is the following statement, which identifies the $\infty$-category $\mathbf{AbChAff}_{k}$ when $k$ is a field. It is the commutative group version of the characterization of affine stacks by means of their homotopy sheaves which is proven in [Toe06].
@@
Proof of the lemma. This is already contained in the proof of [Toe06, Cor. 3.2.7] which we reproduce here for simplicity. Let $A=\mathcal{O}(E)$ be the commutative cosimplicial algebra corresponding to $E$, which is endowed with a $H_{\infty}$-Hopf algebra structure in the sense of [Toe06, Def. 3.1.4], or in other words which is a cogroup in the $\infty$-category of commutative cosimplicial $k$-algebras. As a consequence, $H^{*}(A)$ is a graded commutative Hopf $k$-algebra and therefore each $H^{0}(A)$-module $H^{i}(A)$ is free (because it corresponds to an equivariant quasi-coherent sheaf on the group scheme $\mathbf{Spec}\,H^{0}(A)$), and in particular $A$ is flat as a complex of $H^{0}(A)$-modules. Let $G=\mathbf{Spec}\,H^{0}(A)$ and $p:E\to G$ the natural projection induced by the canonical morphism $H^{0}(A)\to A$. The fiber of $p$ is $\mathbf{Spec}\,A^{\prime}$, where $A^{\prime}=A\otimes_{H^{0}(A)}k$, which, by the flatness of $A$ over $H^{0}(A)$, is reduced $H^{0}(A^{\prime})\simeq k$. By [Toe06, Thm. 2.4.5] the corresponding affine stack $E^{\prime}=\mathbf{Spec}\,A^{\prime}$ is pointed and connected, and thus all the sheaves $\pi_{i}(E^{\prime})$ are representable by unipotent group schemes. The long exact sequence in homotopy therefore implies the lemma, as $\pi_{i}(E)\simeq\pi_{i}(E^{\prime})$ for all $i>0$. $\Box$
The lemma 4.3 as the following important consequence: for any object $E\in\mathbf{AbChAff}_{k}$, the homotopy groups $\pi_{i}(E)$ are all representable by unipotent group schemes, for any $i\geq 0$. Indeed, by the lemma we already know that this is the case for $i>0$. We now consider $E[1]$, and by definition we know that $E[1]$ is again a group object in affine stacks. In particular the lemma implies that $\pi_{1}(E[1])\simeq\pi_{0}(E)$ is representable by a unipotent group scheme.
Conversely, let $E\in\mathbf{AbSt}_{k}$ be an object such that the sheaves $\pi_{i}(E)$ are representable by unipotent group schemes. We have to show that each stack $|E[n]|\in\mathbf{St}_{k}$ is affine. But, the classifying stack $B(|E[n]|)$ is affine because of [Toe06, Thm. 2.4.1], and thus so is $\Omega_{*}B(|E[n]|)\simeq|E[n]|$ as affine stacks are stable by limits. $\Box$
@@
Proof of the corollary. Let $C\subset\mathbf{AbSt}_{k}$ be the smallest sub-$\infty$-category containing $\mathbb{G}_{a}[n]$ and stable by small limits. Clearly $C$ is contained in $\mathbf{AbChAff}_{k}$, and we know that $\mathbf{AbChAff}_{k}$ is stable by limits and that it contains all the objects $\mathbb{G}_{a}[n]$. The converse is a direct consequence of Postnikov decomposition and the proposition. By [Toe20, Prop. B1], Postnikov towers always converge for the fpqc topology (see also [MR22] for a more general statement), and thus any $E\in\mathbf{AbChAff}_{k}$ can be written as $E\simeq\lim_{n}\tau_{\leq n}E$. Moreover, the proposition implies that each $\tau_{n}(E)$ belongs to $\mathbf{AbChAff}_{k}$. Finally, we proceed by induction on $n$ to show that each $\tau_{n}(E)$ belongs to $C$, by means of the cartesian squares
[…]
As $\pi_{n}(E)$ is a unipotent group scheme by proposition 4.2, it is itself contains in $C$, and thus so is $\pi_{n}(E)[n+1]$. $\Box$
@@
Proof. We consider the inclusion of opposite $\infty$-categories. We know that $(\mathbf{ChAff}_{k})^{op}$ is equivalent to the $\infty$-category associated with the combinatorial model category of commutative cosimplicial $k$-algebras by means of the $\mathbf{Spec}$ functor (see [Toe06]). Therefore, $(\mathbf{ChAff}_{k})^{op}$ is a presentable $\infty$-category. Moreover, $(\widehat{\mathbf{AbChAff}_{k}})^{op}$ is generated by colimits by the small set of objects $\mathbb{G}_{a}[n]$. Moreover, when considered as objects in $(\widehat{\mathbf{AbChAff}_{k}})^{op}$, the objects $\mathbb{G}_{a}[n]$ are small, or more precisely $\mathbb{G}_{a}[n]$ are cosmall objects in $\mathbf{AbChAff}_{k}$ as this is implied by the following lemma.
@@
    The object $G$ is truncated in $C$: there is an $n$ such that $Map(E,C)$ is $n$-truncated for all $E\in c$.
@@
where $i\mapsto n(i)$ and $i\mapsto m(i)$ are certain functions. The existence of such a limit decomposition can be seen for instance using MacLane resolutions of abelian group objects as presented in [Bre70, §2]. Because $G$ is $n$-truncated this homotopy limit reduces to a finite homotopy limit, which in turn will commute with filtered colimit. The result follows easily as all objects $G^{m(i)}$ are cocompact in $C$. $\Box$
The lemma implies that $(\widehat{\mathbf{AbChAff}_{k}})^{op}$ is a presentable $\infty$-category It is therefore a formal consequence of Freyd representability theorem (see [NRS20, Thm. 4.1.3]) that the forgetful $\infty$-functor $(\mathbf{AbChAff}_{k})^{op}\rightarrow(\mathbf{ChAff}_{k})^{op}$, which commutes with small colimits, admits a right adjoint. $\Box$
@@
We also remind that there are set theory issues in the situation as the site of affine schemes $Aff_{k}$ is not a small site. As a result, the inclusion $\mathbf{ChAff}_{k}\subset\mathbf{St}_{k}$ does not possess a left adjoint as objects in $\mathbf{St}_{k}$ might not be small. However, we can restrict ourselves to small stacks $\mathbf{St}_{k}^{sm}\subset\mathbf{St}_{k}$, which are all objects that can be written as small colimits of affine schemes. By definition affine stacks are small, as they are spectrum of small commutative cosimplicial $k$-algebras, so we have an inclusion $\mathbf{ChAff}_{k}\hookrightarrow\mathbf{St}_{k}^{sm}$. This last inclusion does have a left adjoint, called the affinization and denote by $F\mapsto(F\otimes k)^{u}$. It is explicitely given by considering $\mathcal{O}(F)$, the commutative cosimplicial $k$-algebra of functions on $F$, which is small as $F$ is assumed to be small, and setting $(F\otimes k)^{u}\simeq\mathbf{Spec}\,\mathcal{O}(F)$.
@@
$\mathbb{H}^{u}(\underline{K})\simeq co\lim_{\Delta(K)}\mathbb{H}^{u}(*),$
and we are thus reduced to compute the affine homology of the point. For this, we recall the existence of the commutative group scheme $\mathbb{H}$ of [Toe20], which can be defined as the spectrum of the Hopf algebra of integral valued polynomials. Equivalently, $\mathbb{H}$ sits inside the group scheme of big Witt vectors $\mathbb{W}$ as the sub-functor of simultaneously fixed points by all Frobenius $F_{p}$. The group scheme $\mathbb{H}$ is the additive part of ring scheme structure, for which the multiplication is induced by the multiplication of Witt vectors.
We claim that $\mathbb{H}^{u}(*)\simeq\mathbb{H}$, considered as a commutative group for the additive structure. First of all, as shown in [Toe20] all the $K(\mathbb{H},n)$ are abelian group objects in affine stacks, and thus $\mathbb{H}$ is indeed an object of $\mathbf{AbChAff}_{k}$. Using the presentation as fixed points of Frobenius on $\mathbb{W}$, it can be checked that $\mathbb{H}$ lies furthermore in $\widehat{\mathbf{AbChAff}_{k}}$ and thus is an affine homology type. In [Toe20] is also known that when $n>0$ $K(\mathbb{H},n)$ is the affinization of the space $K(\mathbb{Z},n)$, and thus the natural morphism $K(\mathbb{Z},n)\to K(\mathbb{H},n)$, which is a morphism of abelian group objects induced by the canonical map $\mathbb{Z}\to\mathbb{H}$, makes $K(\mathbb{H},n)$ into the affinization of $K(\mathbb{Z},n)$. Indeed, for any abelian group object $E$ in $\mathbf{ChAff}_{k}$, the mapping space $Map_{\mathbf{AbSt}_{k}}(K(\mathbb{Z},n),E)$ can be computed by a certain homotopy end
@@
The affinization of $K(\mathbb{Z},n)^{i}\simeq K(\mathbb{Z}^{i},n)$ being $K(\mathbb{H}^{i},n)$ (see [Toe20]), we deduce easily that the natural morphism
@@
is an equivalence for any $E\in\mathbf{AbChAff}_{k}$. In other words, the reduced affine homology of $S^{n}$ is $K(\mathbb{H},n)$, for any $n>0$. But we then have $Map_{\mathbf{AbSt}_{k}}(\mathbb{H},E)\simeq Map_{\mathbf{AbSt}_{k}}(K(\mathbb{H},1),K(E,1))\simeq\Omega_{*}(K(E,1)(k))\simeq E(k)$, and thus $\mathbb{H}$ is indeed the affine homology of * as wanted.
@@
This morphism is not an equivalence in general, but we claim it is so when $k^{\prime}$ is finite and of local complete intersection over $k$. Indeed, by the very definition of affine homology this would follow formally from the fact that the direct image $\infty$-functor $R_{k^{\prime}/k}:\mathbf{AbSt}_{k^{\prime}}\to\mathbf{AbSt}_{k}$, right adjoint to the base change, preserves affine homology types. As these are generated by limits from the $\mathbb{G}_{a}[n]$, it is enough to show that $R_{k^{\prime}/k}(\mathbb{G}_{a}[n])$ is an affine homology type over $k$. But $R_{k^{\prime}/k}(\mathbb{G}_{a}[n])$ is nothing else that $|p_{*}(\mathbb{G}_{a})[n]|$, where $p:\mathbf{Spec}\,k^{\prime}\to\mathbf{Spec}\,k$. Moreover, by assumption on $k^{\prime}/k$ the stack $p_{*}(\mathbb{G}_{a})[n]$ is a linear stack over $k$ associated to the perfect complex $E:=(k^{\prime})^{\vee}[-n]$, which is the dual of $k^{\prime}[n]$. More explicitly, $p_{*}(\mathbb{G}_{a})[n]$ can be written as the spectrum of the cosimplicial commutative $k$-algebras $Sym_{k}(E)$, with $E$ a perfect complex of $k$-modules. These linear stacks, as objects in $\mathbf{AbSt}_{k}$ can be decomposed by considering a projective resolution of $E$, and are easily seen to be expressible by limits of stacks of the form $V[n]$ where $V$ is a vector bundle on $\mathbf{Spec}\,k$, which are retracts of $\mathbb{G}_{a}^{r}[n]$ and thus are affine homology types over $k$.
An important cases of base change is when $k$ is regular, and $k\to k(p)$ is the quotient corresponding to a closed point $p\in\mathbf{Spec}\,k$. In this case base change tells us that $\mathbb{H}^{u}(X/k)$ specializes to $\mathbb{H}^{u}(X_{p}/k(p))$ when pull-backed to the point $p$.
Filtration by dimension. We let $X$ be a $k$-schemes separated and of finite type over $k$. For simplicity we assume that $k$ is noetherian and $X$ is of finite dimension $d$. We can proceed as in definition 3.1 and construct a finite filtration on $\mathbb{H}^{u}(X/k)$
[…]
The associated graded for this filtration, $Gr_{d}\mathbb{H}^{u}(X/k)$, defined considering cofibers in $\widehat{\mathbf{AbChAff}_{k}}$, are affine homology types over $k$ which are related to local cohomology in dimension $d$ as follows. We first introduce affine homology types of finite presentation, as being the objects $E\in\widehat{\mathbf{AbChAff}_{k}}$ being obtained by finite limits and retracts of objects of the form $\mathbb{G}_{a}[n]$. As we have already seen in lemma 4.7 the objects $\mathbb{G}_{a}[n]$ are cocompact in $\mathbf{AbChAff}_{k}$, and thus any affine homology type of finite presentation are cocompact objects amoung affine homology types.

### v3

@@ 3.2. Grothendieck’s construction
Proof of the lemma. Suppose that we are given an extension of commutative pro-algebraic groups ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{G}_{a}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${J_{x}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$ By the universal property of $J_{0}(X)$, construction a section $J_{0}(X)\to H$ is equivalent to give an element $s\in H^{d}_{x}(X,H)$, which projects to the universal element in $H^{d}_{x}(X,J_{x}(X))$. But we have an exact sequence in fppffppf cohomology of $X$ local at $x$ $H^{d}_{x}(X,H)\to H^{d}_{x}(X,J_{x}(X))\to H^{d+1}_{x}(X,\mathbb{G}_{a})=0$, showing that $H^{d}_{x}(X,H)\to H^{d}_{x}(X,J_{x}(X))$ is surjective. Therefore the short exact sequence splits showing that $Ext^{1}_{naive}(J_{x}(X),\mathbb{G}_{a})=0$. Moreover, it is well known, and can be proven using Dieudonné theory, that $Ext^{i}_{naive}(A,\mathbb{G}_{a})=0$ for all $i>1$ and all commutative unipotent pro-algebraic group $A$ over $k$ (see [12] proposition V-1 5.1 and 5.2). $\Box$
We thus know that $\psi_{\mathbb{G}_{a}}^{i}$ is an isomorphism for all $i$. It remains to treat the case of a general commutative unipotent group $G$. Any such group possesses a composition series whose layers $G_{\alpha}$ can be realized in a short exact sequence $0\to G_{\alpha}\to\mathbb{G}_{a}\to\mathbb{G}_{a}$. The associated long exact sequences and the lemma easily imply that $\psi_{G}^{i}$ is also an isomorphism for all $i$ (note that the domains and codomains of $\psi_{G}^{i}$ can be non-zero for $i=1$). $\Box$
@@
    The previous proposition provides a close relation between our algebraic homology $\mathbb{H}^{alg}(X)$ and the original Grothendieck’s construction $J_{*}(X)$. The latter is a truncated version of the former, but both can be used in order to express the fppf cohomology of $X$ with coefficients in a commutative group scheme.
@@
    The extra conditions on $G$, being affine for $i>0$ and unipotent for $i>1$, do not appear in Grothendieck’s original construction. This is because his construction is originaly made using the Zariski topology instead ot he fppf topology, and higher Zariski cohomology groups with coefficients in group schemes vanish outside of the unipotent context.
@@ 4. Partial generalization to more general schemes
In this section we partially extend the results of the previous sections to the more general case of schemes over arbitrary bases. For this, we restrict ourselves to the unipotent part of algebraic homology and use the formalism of commutative groups in affine stacks (in the sense of [32]) as replacement for (pro) commutative group stacks. On the way, we develop the general framework of affine homology, the natural homology theory associated to affine stacks.
@@ 4.1. Affine homology
When $C=\mathbf{St}_{k}$ is the $\infty$-category of stacks over $k$, the $\infty$-category of abelian group objects in $C$ (also called group-complete strictly commutative monoids, see [16, §7]) is naturally equivalent to $\mathbf{AbSt}_{k}$ (already used previously in this work, but in the setting of the fppf topology), and is thus identified with the $\infty$-category of connective complexes of sheaves of abelian groups on the site of affine $k$-schemes with the fpqc topology.
@@
The $\infty$-category of commutative affine group stacks over $k$ is defined to be the full sub-$\infty$-category of $\mathbf{AbSt}_{k}$ consisting of all commutative group stacks $E$ such that for all $n\geq 0$ the underlying stack $|E[n]|\in\mathbf{St}_{k}$ is an affine stack in the sense of [32]. The $\infty$-category of all commutative affine group stacks will be denoted by $\mathbf{AbChAff}_{k}$.
We caution the reader that the notation $\mathbf{AbChAff}_{k}$ can be misleading. Indeed, the $\infty$-category $\mathbf{ChAff}_{k}$ defines another $\infty$-category $Ab(\mathbf{ChAff}_{k})$, of abelian group objects in $\mathbf{ChAff}_{k}$. Recall that this is formally defined as the full sub-$\infty$-category of $Fun^{\infty}(T^{op},\mathbf{ChAff}_{k})$, consisting of all $\infty$-functors sending finite sums in $T$ to finite products in $C$, and where $T$ is the algebraic theory of abelian group (i.e. the opposite of the category of free abelian group of finite rank). The $\infty$-category $\mathbf{AbChAff}_{k}$ is smaller than $Ab(\mathbf{ChAff}_{k})$. Indeed, any affine commutative group scheme $G$ defines an object in $Ab(\mathbf{ChAff}_{k})$. However, we will see below (see Proposition 4.2) that $\mathbb{G}_{m}$ is not an object in $\mathbf{AbChAff}_{k}$. We thus have canonical fully faithful $\infty$-functors $\mathbf{AbChAff}_{k}\hookrightarrow Ab(\mathbf{ChAff}_{k})\hookrightarrow\mathbf{AbSt}_{k}$, but these three $\infty$-categories are all distinct.
@@
An important first observation is the following statement, which identifies the $\infty$-category $\mathbf{AbChAff}_{k}$ when $k$ is a field. It is the commutative group version of the characterization of affine stacks by means of their homotopy sheaves which is proven in [32].
@@
Proof of the lemma. This is already contained in the proof of [32, Cor. 3.2.7] which we reproduce here for simplicity. Let $A=\mathcal{O}(E)$ be the commutative cosimplicial algebra corresponding to $E$, which is endowed with a $H_{\infty}$-Hopf algebra structure in the sense of [32, Def. 3.1.4], or in other words which is a cogroup in the $\infty$-category of commutative cosimplicial $k$-algebras. As a consequence, $H^{*}(A)$ is a graded commutative Hopf $k$-algebra and therefore each $H^{0}(A)$-module $H^{i}(A)$ is free (because it corresponds to an equivariant quasi-coherent sheaf on the group scheme $\mathbf{Spec}\,H^{0}(A)$), and in particular $A$ is flat as a complex of $H^{0}(A)$-modules. Let $G=\mathbf{Spec}\,H^{0}(A)$ and $p:E\to G$ the natural projection induced by the canonical morphism $H^{0}(A)\to A$. The fiber of $p$ is $\mathbf{Spec}\,A^{\prime}$, where $A^{\prime}=A\otimes_{H^{0}(A)}k$, which, by the flatness of $A$ over $H^{0}(A)$, is reduced $H^{0}(A^{\prime})\simeq k$. By [32, Thm. 2.4.5] the corresponding affine stack $E^{\prime}=\mathbf{Spec}\,A^{\prime}$ is pointed and connected, and thus all the sheaves $\pi_{i}(E^{\prime})$ are representable by unipotent group schemes. The long exact sequence in homotopy therefore implies the lemma, as $\pi_{i}(E)\simeq\pi_{i}(E^{\prime})$ for all $i>0$. $\Box$
The lemma 4.3 has the following important consequence: for any object $E\in\mathbf{AbChAff}_{k}$, the homotopy groups $\pi_{i}(E)$ are all representable by unipotent group schemes, for any $i\geq 0$. Indeed, by the lemma we already know that this is the case for $i>0$. We now consider $E[1]$, and by definition we know that $E[1]$ is again a group object in affine stacks. In particular the lemma implies that $\pi_{1}(E[1])\simeq\pi_{0}(E)$ is representable by a unipotent group scheme.
Conversely, let $E\in\mathbf{AbSt}_{k}$ be an object such that the sheaves $\pi_{i}(E)$ are representable by unipotent group schemes. We have to show that each stack $|E[n]|\in\mathbf{St}_{k}$ is affine. But, the classifying stack $B(|E[n]|)$ is affine because of [32, Thm. 2.4.1], and thus so is $\Omega_{*}B(|E[n]|)\simeq|E[n]|$ as affine stacks are stable by limits. $\Box$
@@
Proof of the corollary. Let $C\subset\mathbf{AbSt}_{k}$ be the smallest sub-$\infty$-category containing $\mathbb{G}_{a}[n]$ and stable by small limits. Clearly $C$ is contained in $\mathbf{AbChAff}_{k}$, and we know that $\mathbf{AbChAff}_{k}$ is stable by limits and that it contains all the objects $\mathbb{G}_{a}[n]$. The converse is a direct consequence of Postnikov decomposition and the proposition. By [34, Prop. B1], Postnikov towers always converge for the fpqc topology (see also [25] for a more general statement), and thus any $E\in\mathbf{AbChAff}_{k}$ can be written as $E\simeq\lim_{n}\tau_{\leq n}E$. Moreover, the proposition implies that each $\tau_{\leq n}(E)$ belongs to $\mathbf{AbChAff}_{k}$. Finally, we proceed by induction on $n$ to show that each $\tau_{\leq n}(E)$ belongs to $C$, by means of the cartesian squares
${\tau_{\leq n}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\tau_{\leq n-1}(E)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\pi_{\leq n}(E)[n+1].}$
As $\pi_{\leq n}(E)$ is a unipotent group scheme by the proposition 4.2, it is itself contains in $C$, and thus so is $\pi_{\leq n}(E)[n+1]$. $\Box$
@@
Proof. We consider the inclusion of opposite $\infty$-categories. We know that $(\mathbf{ChAff}_{k})^{op}$ is equivalent to the $\infty$-category associated with the combinatorial model category of commutative cosimplicial $k$-algebras by means of the $\mathbf{Spec}$ functor (see [32]). Therefore, $(\mathbf{ChAff}_{k})^{op}$ is a presentable $\infty$-category. Moreover, $(\widehat{\mathbf{AbChAff}_{k}})^{op}$ is generated by colimits by the small set of objects $\mathbb{G}_{a}[n]$. Moreover, when considered as objects in $(\widehat{\mathbf{AbChAff}_{k}})^{op}$, the objects $\mathbb{G}_{a}[n]$ are small, or more precisely $\mathbb{G}_{a}[n]$ are cosmall objects in $\mathbf{AbChAff}_{k}$ as this is implied by the following lemma.
@@
    The object $G$ is truncated in $C$: there is an $n$ such that $Map(E,G)$ is $n$-truncated for all $E\in C$.
@@
where $i\mapsto n(i)$ and $i\mapsto m(i)$ are certain functions. The existence of such a limit decomposition can be seen for instance using MacLane resolutions of abelian group objects as presented in [5, §2]. Because $G$ is $n$-truncated this homotopy limit reduces to a finite homotopy limit, which in turn will commute with filtered colimit. The result follows easily as all objects $G^{m(i)}$ are cocompact in $C$. $\Box$
The lemma implies that $(\widehat{\mathbf{AbChAff}_{k}})^{op}$ is a presentable $\infty$-category It is therefore a formal consequence of Freyd representability theorem (see [26, Thm. 4.1.3]) that the forgetful $\infty$-functor $(\widehat{\mathbf{AbChAff}_{k}})^{op}\rightarrow(\mathbf{ChAff}_{k})^{op}$, which commutes with small colimits, admits a right adjoint. $\Box$
@@
We also remind that there are set theory issues in the situation as the site of affine schemes $Aff_{k}$ is not a small site. As a result, the inclusion $\mathbf{ChAff}_{k}\subset\mathbf{St}_{k}$ does not possess a left adjoint as objects in $\mathbf{St}_{k}$ might not be small. However, we can restrict ourselves to small stacks $\mathbf{St}_{k}^{sm}\subset\mathbf{St}_{k}$, which are all objects that can be written as small colimits of affine schemes. By definition affine stacks are small, as they are spectrum of small commutative cosimplicial $k$-algebras, so we have an inclusion $\mathbf{ChAff}_{k}\hookrightarrow\mathbf{St}_{k}^{sm}$. This last inclusion does have a left adjoint, called the affinization and denoted by $F\mapsto(F\otimes k)^{u}$. It is explicitely given by considering $\mathcal{O}(F)$, the commutative cosimplicial $k$-algebra of functions on $F$, which is small as $F$ is assumed to be small, and setting $(F\otimes k)^{u}\simeq\mathbf{Spec}\,\mathcal{O}(F)$.
@@
$\mathbb{H}^{u}(\underline{K})\simeq\underset{\Delta(K)}{colim}\mathbb{H}^{u}(*),$
and we are thus reduced to compute the affine homology of the point. For this, we recall the existence of the commutative group scheme $\mathbb{H}$ of [34], which can be defined as the spectrum of the Hopf algebra of integral valued polynomials. Equivalently, $\mathbb{H}$ sits inside the group scheme of big Witt vectors $\mathbb{W}$ as the sub-functor of simultaneously fixed points by all Frobenius $F_{p}$. The group scheme $\mathbb{H}$ is the additive part of ring scheme structure, for which the multiplication is induced by the multiplication of Witt vectors.
We claim that $\mathbb{H}^{u}(*)\simeq\mathbb{H}$, considered as a commutative group for the additive structure. First of all, as shown in [34] all the $K(\mathbb{H},n)$ are abelian group objects in affine stacks, and thus $\mathbb{H}$ is indeed an object of $\mathbf{AbChAff}_{k}$. Using the presentation as fixed points of Frobenius on $\mathbb{W}$, it can be checked that $\mathbb{H}$ lies furthermore in $\widehat{\mathbf{AbChAff}_{k}}$ and thus is an affine homology type. In [34] is also known that when $n>0$ $K(\mathbb{H},n)$ is the affinization of the space $K(\mathbb{Z},n)$, and thus the natural morphism $K(\mathbb{Z},n)\to K(\mathbb{H},n)$, which is a morphism of abelian group objects induced by the canonical map $\mathbb{Z}\to\mathbb{H}$, makes $K(\mathbb{H},n)$ into the affinization of $K(\mathbb{Z},n)$. Indeed, for any abelian group object $E$ in $\mathbf{ChAff}_{k}$, the mapping space $Map_{\mathbf{AbSt}_{k}}(K(\mathbb{Z},n),E)$ can be computed by a certain homotopy end
@@
The affinization of $K(\mathbb{Z},n)^{i}\simeq K(\mathbb{Z}^{i},n)$ being $K(\mathbb{H}^{i},n)$ (see [34]), we deduce easily that the natural morphism
@@
is an equivalence for any $E\in\mathbf{AbChAff}_{k}$. In other words, the reduced affine homology of $S^{n}$ is $K(\mathbb{H},n)$, for any $n>0$. But we then have $Map_{\mathbf{AbChAff}_{k}}(\mathbb{H},E)\simeq Map_{\mathbf{AbChAff}_{k}}(K(\mathbb{H},1),K(E,1))\simeq\Omega_{*}(K(E,1)(k))\simeq E(k)$, and thus $\mathbb{H}$ is indeed the affine homology of * as wanted.
@@
This morphism is not an equivalence in general, but we claim it is so when $k^{\prime}$ is finite and of local complete intersection over $k$. Indeed, by the very definition of affine homology this would follow formally from the fact that the direct image $\infty$-functor $R_{k^{\prime}/k}:\mathbf{AbSt}_{k^{\prime}}\to\mathbf{AbSt}_{k}$, right adjoint to the base change, preserves affine homology types. As these are generated by limits from the $\mathbb{G}_{a}[n]$, it is enough to show that $R_{k^{\prime}/k}(\mathbb{G}_{a}[n])$ is an affine homology type over $k$. But $R_{k^{\prime}/k}(\mathbb{G}_{a}[n])$ is nothing else than $|p_{*}(\mathbb{G}_{a})[n]|$, where $p:\mathbf{Spec}\,k^{\prime}\to\mathbf{Spec}\,k$. Moreover, by assumption on $k^{\prime}/k$ the stack $p_{*}(\mathbb{G}_{a})[n]$ is a linear stack over $k$ associated to the perfect complex $E:=(k^{\prime})^{\vee}[-n]$, which is the dual of $k^{\prime}[n]$. More explicitly, $p_{*}(\mathbb{G}_{a})[n]$ can be written as the spectrum of the cosimplicial commutative $k$-algebras $Sym_{k}(E)$, with $E$ a perfect complex of $k$-modules. These linear stacks, as objects in $\mathbf{AbSt}_{k}$ can be decomposed by considering a projective resolution of $E$, and are easily seen to be expressible by limits of stacks of the form $V[n]$ where $V$ is a vector bundle on $\mathbf{Spec}\,k$, which are retracts of $\mathbb{G}_{a}^{r}[n]$ and thus are affine homology types over $k$.
An important case of base change is when $k$ is regular, and $k\to k(p)$ is the quotient corresponding to a closed point $p\in\mathbf{Spec}\,k$. In this case base change tells us that $\mathbb{H}^{u}(X/k)$ specializes to $\mathbb{H}^{u}(X_{p}/k(p))$ when pull-backed to the point $p$.
Filtration by dimension. We let $X$ be a $k$-scheme separated and of finite type over $k$. For simplicity we assume that $k$ is noetherian and $X$ is of finite dimension $d$. We can proceed as in the definition 3.1 and construct a finite filtration on $\mathbb{H}^{u}(X/k)$
${0=F_{d+1}\mathbb{H}^{u}(X/k)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{d}\mathbb{H}^{u}(X/k)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\dots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{1}\mathbb{H}^{u}(X/k)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${F_{0}\mathbb{H}^{u}(X/k)=\mathbb{H}^{u}(X/k).}$
The associated graded for this filtration, $Gr_{d}\mathbb{H}^{u}(X/k)$, defined considering cofibers in $\widehat{\mathbf{AbChAff}_{k}}$, are affine homology types over $k$ which are related to local cohomology in dimension $d$ as follows. We first introduce affine homology types of finite presentation, as being the objects $E\in\widehat{\mathbf{AbChAff}_{k}}$ being obtained by finite limits and retracts of objects of the form $\mathbb{G}_{a}[n]$. As we have already seen in the lemma 4.7 the objects $\mathbb{G}_{a}[n]$ are cocompact in $\mathbf{AbChAff}_{k}$, and thus any affine homology type of finite presentation are cocompact objects among affine homology types.